# Tran Wang Ge Liu Yamagishi

### v1

@@ 3.4.2 Matched data domain but different synthetic methods
When evaluating across the test sets in the same audiobook domain, the performance of fine-tuned Whisper and the ResNet degraded to different degrees.
On E.Voc, the fine-tuned Whisper and the ResNet trained from scratch using Ft.TTS obtained higher FAR and FRR than their counterparts using Ft.Voc. For analysis, we calculated the FAR and FRR values evaluated after grouping the words based on their durations. Figure 2 (a) plots the results of the Whisper fine-tuned on Ft.Voc (black profiles) or Ft.TTS (red profiles). While the FARs of the two fine-tuned Whispers are close, the FRR when using Ft.TTS (red dashed line) is around or higher than 80%. In contrast, the FRR in the case of using Ft.Voc (black dashed line) is close to 0%.
To further investigate the above result, we analyzed the errors across the languages using E.Voc. The results presented in Table 3 show that using Ft.TTS led to FAR and FRR above 40% on the English data. On other languages, it obtained FARs similar to those in the case of using Ft.Voc but FRRs close to 90%. This suggests that the increased FRR when training on Ft.TTS but testing on E.Voc is due to the unseen languages.
On E.TTS, the results in Table 2 show that the Whisper fine-tuned on Ft.Voc obtained higher FAR and FRR than its counterparts using Ft.TTS. Particularly, its FAR was as high as 76%. As analysis, we plot the FARs and FRRs against the word duration in Figure 2 (b) and observe that the Whisper fine-tuned on Ft.Voc obtained high FAR (black solid line) regardless of the word duration. The high FAR values suggest that the Whisper fine-tuned on the vocoded training data struggles to spot words created by the TTS systems. Using more diverse vocoders may improve the performance, but the exploration is left to future work.
Note that the FRR of the Whisper fine-tuned on Ft.Voc increases when the word is longer (black dashed line in Figure 2 (b)). The same trend is observed on the Whisper fine-tuned using Ft.TTS (red dashed line). One hypothesis is that the fine-tuned Whisper seems to flag only if it does not find any artifact within a word. The longer a word is, the more likely it contains patterns similar to artifacts.
@@ 3.5 Results on out-of-domain test data
Given the results above, it is not surprising that the performance of the fine-tuned Whisper and the ResNet also degraded on the out-of-domain test data. As Table 4 shows, on both E.AV1M and E.PE, we observe higher WER when the Whisper is fine-tuned using either vocoded or TTS-synthesized data. In terms of synthetic word detection, the detection performance is overall unstable. For example, in the case of training using Ft.Voc and testing on E.PE, both models obtained low FRRs ( 10%) but very high FARs (  80%). In short, current out-of-domain synthetic word detection is not yet sufficiently reliable.
Note that, in the case of using Ft.TTS, comparison between the results on E.TTS and E.AV1M shows that both the fine-tuned Whisper and ResNet performed worse on E.AV1M, even though the synthesizers (YourTTS and VITS) are covered by the training data. Hence, to improve the out-of-domain performance, understanding the differences across datasets may be more important than complicating the model architecture.
Table 3: Decomposed results over languages, using the pre-trained Whisper (left) or fine-tuned on Ft.Voc (right).
@@
Table 4: Results on test sets manipulated by DNN-based speech synthesis and editing systems. Within each column, a value with a darker color indicates a worse performance.
@@ 4 Conclusion
We proposed a method for fine-tuning Whisper to detect synthetic words while preserving its speech-to-text transcription capability. The proposed method requires only augmenting the text token sequences with tokens that mark synthetic words, with no change to the algorithm or model architecture. This integrates the detection task into the next-token prediction task. In the experiment wherein the training and test data matched, the fine-tuned Whisper, while doing transcription, detected the synthetic words with a performance comparable to a dedicated ResNet-based detection model. However, results on multiple test sets also demonstrate that the detection performance degraded when there is a data domain mismatch between the training and test sets. A better understanding of the cross-domain differences is needed for future work. Using vocoded training data did not alleviate the degradation when the synthetic words are from a different domain or unseen synthesizers. Future work will investigate using vocoded data from diverse synthesizers and domains.

### v2

@@ 3.4.2 Matched data domain but different synthetic methods
When evaluating across the test sets in the same audiobook domain, the performance of fine-tuned Whisper and ResNet degraded to different degrees.
On E.Voc, the fine-tuned Whisper and ResNet trained from scratch using Ft.TTS obtained higher FARs and FRRs than their counterparts using Ft.Voc. For analysis, we calculated the FARs and FRRs after grouping the words on the basis of their durations. Figure 2 (a) plots the results of Whisper fine-tuned on Ft.Voc (black profiles) or Ft.TTS (red profiles). While the FARs of the two fine-tuned Whispers are close, the FRR when using Ft.TTS (red dashed line) is around or higher than 80%. In contrast, the FRR when using Ft.Voc (black dashed line) is close to 0%.
To further investigate the above results, we analyzed the errors across the languages using E.Voc. The results presented in Table 3 indicate that using Ft.TTS led to FAR and FRR above 40% on the English data. On other languages, it obtained FARs similar to those when using Ft.Voc but FRRs close to 90%. This suggests that the increased FRR when training on Ft.TTS but testing on E.Voc is due to the unseen languages.
On E.TTS, the results in Table 2 indicate that Whisper fine-tuned on Ft.Voc obtained higher FAR and FRR than its counterparts using Ft.TTS. Particularly, its FAR was as high as 76.16%. As analysis, we plot the FARs and FRRs against the word duration in Figure 2 (b) and observe that Whisper fine-tuned on Ft.Voc obtained high FARs (black solid line) regardless of the word duration. A high FAR suggests that Whisper fine-tuned on the vocoded training data struggles to spot words created using the TTS systems. Using more diverse vocoders may improve performance, but exploration is left to future work.
Note that the FRR of Whisper fine-tuned on Ft.Voc increases when the word is longer (black dashed line in Figure 2 (b)). The same trend is observed on Whisper fine-tuned using Ft.TTS (red dashed line). One hypothesis is that the fine-tuned Whisper seems to flag only if it does not find any artifact within a word. The longer a word is, the more likely it contains patterns similar to artifacts.
@@ 3.5 Results on out-of-domain test data
Given the above results, it is not surprising that the performance of the fine-tuned Whisper and ResNet also degraded on the out-of-domain test data. As Table 4 shows, on both E.AV1M and E.PE, we observe higher WER when Whisper is fine-tuned using either vocoded or TTS-synthesized data. In terms of synthetic word detection, the detection performance is overall unstable. For example, when training using Ft.Voc and testing on E.PE, both models obtained low FRRs ( 10%) but very high FARs (  80%). In short, current out-of-domain synthetic word detection is not yet sufficiently reliable.
Note that, when using Ft.TTS, comparison between the results on E.TTS and E.AV1M shows that both the fine-tuned Whisper and ResNet performed worse on E.AV1M, even though the synthesizers (YourTTS and VITS) were covered by the training data. Hence, to improve the out-of-domain performance, understanding the differences across datasets may be more important than complicating the model architecture.
Table 3: Decomposed results over languages using pre-trained Whisper (left) or fine-tuned on Ft.Voc (right).
@@
Table 4: Results on test sets manipulated using DNN-based speech synthesis and editing systems. Within each column, value with darker shading indicates worse performance.
@@ 4 Conclusion
We proposed a method for fine-tuning Whisper to detect synthetic words while preserving its speech-to-text transcription capability. The proposed method requires only augmenting the text token sequences with tokens that mark synthetic words, with no change to the algorithm or model architecture. This integrates the detection task into the next-token prediction task. In the experiment in which the training and test data matched, the fine-tuned Whisper, while executing transcription, detected the synthetic words with a performance comparable to a dedicated ResNet-based detection model. However, results on multiple test sets also indicate that the detection performance degraded when there was a data-domain mismatch between the training and test sets. A better understanding of the cross-domain differences is needed for future work. Using vocoded training data did not mitigate the degradation when the synthetic words were from a different domain or unseen synthesizers. Future work will investigate using vocoded data from diverse synthesizers and domains.